OpenStudy (anonymous):

Mr. Hauseman has 17 students in his class, three of whom are freshmen, and the rest are from other classes. He is going to draw two students randomly to be partners. He calculates the probability of drawing a freshman and then a junior to be 9/136. How many juniors must be in the class?

OpenStudy (anonymous):

Suppose that order does not matter. (Freshman + junior = junior + Freshman) Let a be number of juniors in class P(drawing junior and freshman) = 3/17 x a/16 9/136 = 3a / 272 9(272) = (3a)(136) 2448 = 408a a = 6

OpenStudy (anonymous):

it is 6

OpenStudy (anonymous):

A standard deck of playing cards contains 52 cards, equally divided among four suits (hearts, diamonds, clubs, and spades). Each suit has the cards 2 through 10, as well as a jack, a queen, a king, and an ace. If the 3 of spades is drawn from a standard deck and is not replaced, what is the probability that the next card drawn is a spade OR a king?